# LLM Roleplay: Simulating Human-Chatbot Interaction

### v1

@@ Abstract
The development of chatbots requires collecting a large number of human-chatbot dialogues to reflect the breadth of users’ sociodemographic backgrounds and conversational goals. However, the resource requirements to conduct the respective user studies can be prohibitively high and often only allow for a narrow analysis of specific dialogue goals and participant demographics. In this paper, we propose LLM-Roleplay: a goal-oriented, persona-based method to automatically generate diverse multi-turn dialogues simulating human-chatbot interaction. LLM-Roleplay can be applied to generate dialogues with any type of chatbot and uses large language models (LLMs) to play the role of textually described personas. To validate our method we collect natural human-chatbot dialogues from different sociodemographic groups and conduct a human evaluation to compare real human-chatbot dialogues with our generated dialogues. We compare the abilities of state-of-the-art LLMs in embodying personas and holding a conversation and find that our method can simulate human-chatbot dialogues with a high indistinguishability rate.111https://github.com/UKPLab/llm-roleplay
T1zi4lmr
@@ 1 Introduction
Collecting human-chatbot dialogues requires recruiting and managing a large number of human annotators, which can pose prohibitive obstacles to researchers who aim to develop conversational AI agents (i.e., chatbots). To circumvent the latter and the limitations of publicly available data, numerous methods employing chatbots to generate dialogues have been introduced lately Xu et al. (2023b); Kim et al. (2023); Zhu et al. (2023); Ding et al. (2023); Zhao et al. (2024b). These methods have the potential to generate synthetic data more quickly and cost-effectively, offering a level of quality and variety that can closely match what human annotators can achieve Zhang et al. (2024).
@@
Recently, several studies have leveraged such synthetic data generated through distillation and self-improvement techniques to enhance the capabilities of pre-trained large language models (LLMs) Zhang et al. (2024). By employing instruction tuning and fine-tuning techniques, these efforts aim to ensure that LLMs deliver useful and safe responses that adhere to provided instructions Xu et al. (2023a); Taori et al. (2023); Mukherjee et al. (2023); Li et al. (2023b); Peng et al. (2023); Almazrouei et al. (2023).
@@
However, today’s dialogue generation methods have two critical limitations. First, the set of possible user responses for a given chatbot utterance is large while the existing datasets typically only cover a single or few annotated user responses. The previously generated datasets are constrained by the number of dialogues and turns, making it difficult to extend them to novel domains or conversational goals. Second, the existing methods rarely account for subjective response behavior and implicitly assume an average user, which can create a gap in representation Rottger et al. (2022) and poses the danger of evaluating and ultimately improving chatbots for a specific sociodemographic group while ignoring others.
@@
The quality of existing datasets is primarily evaluated with the performance in a supervised fine-tuning (SFT) setup, which is influenced by various properties of the generated dataset. For example, Chen et al. (2023) argue that the initial prompt plays a crucial role in the quality of the generated dialogue. On the other hand, Zhao et al. (2024a) show that the length of the response is the key factor and with significantly fewer samples outperforms methods that solely focus on quality. Meanwhile, Shen (2024) emphasizes the importance of mimicking human-style interactions in the dialogues.
@@
To mitigate the shortcomings of existing dialogue generation methods and to account for the desiderate identified in prior work for SFT, we introduce LLM Roleplay: a goal-oriented, persona-based, plug-and-play method to generate diverse, multi-turn, and long dialogues simulating human-chatbot interactions using LLMs. Our method is the first to instruct an LLM (inquirer) to mimic a given persona and prompt a chatbot (responder) to achieve a given conversational goal and thereby elicit realistic human-AI interactions with a specific LLM. Figure 1 shows an exemplary application of our method.
@@
In this work, we address the following two research questions: (i) To what extent can we simulate real human-chatbot dialogues using LLM-chatbot dialogues? (ii) How do various LLMs compare as inquirers within this setup? To answer those research questions, we conduct two user studies. First, we collect real human-chatbot dialogues by asking participants to reach various conversational goals and link the collected dialogues with participants’ sociodemographic background. Next, we use LLM-Roleplay to simulate dialogues with the same set of personas and goals. Second, we conduct a human evaluation with another pool of participants to assess how well the generated dialogues mimic the collected ones. Specifically, we present two dialogues: one collected from the previous study and one simulated using LLM Roleplay, using the same persona and the conversation goal. We then ask the participants to try to identify the simulated one similar to a Turing test Turing (1950).
@@
Overall, this paper contributes: (i) A novel method to simulate human-chatbot dialogues with an arbitrary set of personas and conversational goals, (ii) a human validation study that confirms our method’s potential to simulate human-chatbot interaction with high resemblance to real dialogues, (iii) a dataset of goal-oriented human-chatbot and model-chatbot dialogues using our hand-crafted multi-hop goals involving four state-of-the-art LLMs, (iv) an in-depth comparison of open-source and proprietary LLMs regarding their ability to embody persona and hold conversation, (v) an open-source implementation of our plug-and-play method that can be directly applied to simulate dialogues with any combination of models and chatbot systems using any set of conversational goals and personas.
@@ 2 Large Language Model Roleplay
At a high level, LLM-Roleplay consists of three main steps: (i) Initial conditioning of the inquirer model with persona-specific text alongside the goal description. (ii) Subsequently, the inquirer’s output is provided to the responder model. (iii) Lastly, the output of the responder is returned to the inquirer by asking it to either output a follow-up question or a pre-defined token to terminate the dialogue.
We begin with assembling prompting templates for both of the subjects. We create a system prompt template (SYS_I) for the inquirer LLM. This template prompts the LLM to embody the given persona, provide a prompt to address the designated goal, and output the specified termination token if it considers the goal accomplished. As for the responder, we use a default system prompt (SYS_R) to promote it to be a helpful and honest assistant. Additionally, we develop a response forwarder template (INTER_I) for the inquirer. This template prompts the inquirer to assess the conclusiveness of the answer of the responder, determining whether to output a subsequent question or the termination token. We provide the system and forwarder prompt templates in Table 5 in Appendix B. The algorithm for the LLM Roleplay is shown in Algorithm 1.
@@ 3.1 Persona-Specific Dialogue Collection
Subsequently, we generate dialogues utilizing our LLM Roleplay method with a set of three state-of-the-art LLMs and one proprietary conversational agent as inquirers (): llama-2-13B-Chat (Llama-2) Touvron et al. (2023), Mixtral-8x7B-Instruct-v0.1 (Mixtral) Jiang et al. (2024), vicuna-13b-v1.5-16k (Vicuna) Peng et al. (2023), and GPT4 OpenAI et al. (2024). See sample natural and generated dialogues using Mixtral inquirer and Llama-2 responder from 14 to 18 in Appendix E.
Eventually, the dialogue collection and generation results in the creation of 200 natural human-chatbot () and 800 (with 4 inquirers) simulated () dialogue pairs.
@@ 3.2 Failure Cases
Since the outputs of the LLMs are free-form, applying deterministic functions to their output proves challenging, resulting in a theoretically infinite number of turns. Therefore, there is a need for a set limit on the number of turns: . Nevertheless, we try to capture the algorithm failure cases to have an automatic assessment of the inquirer LLM failure cases. We expect the inquirer LLM to provide the intended prompt enclosed in double quotes as we explicitly request this in our instructions. See a sample expected output and the failure cases examples in Table 7 in Appendix C.
All of the LLMs in our experiments face the following issues.
@@
| Model | Detection Prob.* | Utterance Num.* | Duration |
| Llama-2 | B | C | A |
| Mixtral | A | B | A |
| Vicuna | C | B | A |
| GPT4 | AB | A | A |
@@
In most cases, GPT4 outperforms other models, exhibiting lower occurrences of failure cases: 0.17% for responses without prompts, 0.03% for incoherent responses, and 5.48% for self-replies (Table 1). However, it struggles with providing a single prompt, failing 39.21% of the time. On the contrary, Vicuna excels in generating single prompts, achieving a failure rate of 6.08%. Llama-2 seems to receive fewer incoherent outputs from the responder (also Llama-2) with 0.56%. Mixtral’s performance lies somewhere in between, with more balanced results across different metrics. Refere to Figure 8 in Appendix B for a radar chart. Refer to the detailed data in Table 6 in Appendix B.
@@ 3.3 Human-Evaluation
To further investigate how various instruction-tuned LLMs behave as inquirers within this setup we analyze the simulated dialogue detection probability, the detection utterance number, and the duration participants spent, using generalized linear mixed models (GLMMs), with the choice of LLM as the independent variable. We additionally include random effects to account for potential confounding effects of individual participant’s detection abilities and users from Section 3.1.
We find significant effects of the choice of the model on the detection probability and the utterance position at which the participants formed their decision. We summarize the results of our statistical analyses in Table 3 using CLD codings (Piepho, 2004) and discuss our key findings in the following.
@@
First, we fit a binomial model (logit link) to predict the detection probability depending on the model. Concretely, we estimate a GLMM specified by: . We observe a significant effect of the choice of the model on the detection probability (=21.49, ). A post hoc Wald comparison of the contrasts for model types revealed significant differences between Vicuna and all other models (Vicuna being most likely to be detected), and Mixtral and LLama-2 (Mixtral being significantly less likely to be detected). We do not find a significant difference between GPT4 and Mixtral.
@@
Overall, our findings suggest that Mixtral, an open-source LLM, demonstrates better performance in embodying a specified persona and simulating human-chatbot interactions compared to GPT4 in terms of detection probabilities. Nonetheless, when considering the detection of simulated dialogues based on the detection utterance number, GPT4 shows better performance which is likely to be attributable to GPT4 generating dialogues with more utterances.
@@ 4 Related Work
Approximation of the true distribution of human-chatbot dialogues is challenging and a significant number of diverse human annotators are needed. Incorporating participants from different sociodemographic profiles addressing a given conversational goal can substantially enhance the richness of the dataset. However, the associated time and resource requirements render the respectively needed large-scale studies infeasible for many researchers. Alternatively, a less resource-intensive and faster method for collecting human-chatbot conversations is to generate them using large language models (LLMs) Zhang et al. (2024). This method has been utilized in various setups for dialogue generation, such as human-human Adiwardana et al. (2020); Kim et al. (2023); Chen et al. (2023); Li et al. (2023a), teacher-student Macina et al. (2023), and patient-physician Wang et al. (2023).
@@
Large language models (LLMs) have demonstrated distinct behaviors and personas Andreas (2022); Wolf et al. (2024). According to Andreas (2022), LLMs can interpret behaviors from text prompts, influencing the content they generate. Additionally, Wolf et al. (2024) argue that LLMs function as mixture decompositions, where prompts can shift the balance of these components, thus triggering persona-specific responses. Furthermore, Beck et al. (2024) investigate the effects of sociodemographic prompts on model responses, finding that while such prompts can be beneficial in certain settings, they produce varied effects across different models.
Building on these findings, our work introduces a novel method that is goal-oriented and persona-centric for generating diverse and multi-turn dialogues. This method allows for the simulation of dialogues across various combinations of conversational goals, personas, and LLMs, facilitating the creation of numerous simulated dialogues without theoretical limits.
@@ 5 Discussion and Future Work
In this work, we propose our novel LLM Roleplay method: an automatic, model-agnostic approach to elicit multi-turn, goal-oriented, persona-based simulated human-chatbot dialogues. We develop and validate our method along with two user studies with a total of 40 participants. Our findings indicate that our method can produce dialogues that, in up to 44% (where 50% corresponds to perfect indistinguishability), are indistinguishable from real human-chatbot dialogues and that the respective dialogues cover a higher number of turns than previously proposed datasets.
Building on the findings of this work, several avenues for future research warrant exploration. First, our method can be used to generate user-specific conversational datasets for more targeted alignment and domain-adaption (e.g., using RLHF Christiano et al. (2017)). Second, our method can improve dialogue evaluation by providing high amounts of realistic conversational data.
While this paper presents initial findings on LLM Roleplay and provides the first evidence that our method can approximate human-chatbot dialogues, the validity of our method regarding its sociodemographic representativeness remains to be assessed and advanced. We release our collected dialogue dataset, the code to apply our method, as well as our synthesized dialogues, to facilitate future work towards this promising field of research.
@@ 6 Conclusion
We present our novel LLM Roleplay method for simulating human-chatbot interaction. In a series of two user studies, we collect real human-chatbot dialogues and demonstrate that LLM Roleplay can generate diverse multi-turn conversations that approximate natural human-chatbot dialogues with a high level of indistinguishability. Our findings highlight the potential of LLMs in embodying specified personas and simulating human-chatbot interactions to synthesize realistic dialogues that create new opportunities for real-time model evaluation and training data generation for model fine-tuning.

### v2

@@ Abstract
The development of chatbots requires collecting a large number of human-chatbot dialogues to reflect the breadth of users’ sociodemographic backgrounds and conversational goals. However, the resource requirements to conduct the respective user studies can be prohibitively high and often only allow for a narrow analysis of specific dialogue goals and participant demographics. In this paper, we propose LLM Roleplay: a goal-oriented, persona-based method to automatically generate diverse multi-turn dialogues simulating human-chatbot interaction. LLM Roleplay can be applied to generate dialogues with any type of chatbot and uses large language models (LLMs) to play the role of textually described personas. To validate our method, we collect natural human-chatbot dialogues from different sociodemographic groups and conduct a user study to compare these with our generated dialogues. We evaluate the capabilities of state-of-the-art LLMs in maintaining a conversation during their embodiment of a specific persona and find that our method can simulate human-chatbot dialogues with a high indistinguishability rate.111https://github.com/UKPLab/llm-roleplay
@@ 1 Introduction
Collecting human-chatbot dialogues requires recruiting and managing a large number of human annotators, which can pose prohibitive obstacles to researchers who aim to develop conversational AI agents (i.e., chatbots). To circumvent the latter and the limitations of publicly available data, numerous methods employing chatbots to generate dialogues have been introduced lately Xu et al. (2023b); Kim et al. (2023); Zhu et al. (2023); Ding et al. (2023); Svikhnushina and Pu (2023); Zhao et al. (2024b). These methods can generate dialogues much faster and more cost-effectively while still approximating the quality and variety of human annotators. Zhang et al. (2024).
@@
Recently, several studies have leveraged such synthetic data generated through distillation and self-improvement techniques to enhance the capabilities of large language models (LLMs) Zhang et al. (2024). By employing instruction tuning and fine-tuning techniques, these efforts aim to ensure that LLMs generate useful and safe responses that adhere to provided instructions Xu et al. (2023a); Taori et al. (2023); Mukherjee et al. (2023); Li et al. (2023b); Peng et al. (2023); Almazrouei et al. (2023).
@@
However, current dialogue generation methods have two critical limitations. First, the set of possible user responses for a given chatbot utterance is large while the existing datasets typically only cover a single or few annotated user responses. Previously generated datasets are constrained by the number of dialogues and turns, making it difficult to extend them to novel domains or conversational goals. Second, existing methods rarely account for subjective response behavior and implicitly assume an average user. This can create a gap in representation Rottger et al. (2022) and pose the risk of evaluating and improving chatbots for a specific sociodemographic group while neglecting others.
@@
The quality of existing datasets is primarily assessed through model performance in a supervised fine-tuning (SFT) setup, which is influenced by various properties of the generated dataset. For example, Chen et al. (2023) argue that the initial prompt plays a crucial role in the quality of the generated dialogue. On the other hand, Zhao et al. (2024a) demonstrate that the length of the response is a key factor, and with significantly fewer samples, it outperforms methods that focus solely on quality. Meanwhile, Shen (2024) emphasizes the importance of mimicking human-style interactions in the dialogues.
@@
To mitigate the shortcomings of existing dialogue generation methods and respond to the needs identified in prior work for SFT, we introduce LLM Roleplay: a goal-oriented, persona-based, plug-and-play method designed to generate diverse, multi-turn, long dialogues that simulate human-chatbot interactions using LLMs. Our method is the first to instruct an LLM (inquirer) to adopt a specific persona and prompt a chatbot (responder) to achieve a given conversational goal, thereby eliciting realistic human-AI interactions with a particular LLM. Figure 1 illustrates an exemplary application of our method.
@@
In this work, we address the following two research questions: (i) To what extent can we simulate real human-chatbot dialogues using LLM-chatbot dialogues? (ii) How do various LLMs perform as inquirers within this setup? To answer those research questions, we conduct two user studies. First, we collect real human-chatbot dialogues by asking participants to reach various conversational goals and link the collected dialogues with participants’ sociodemographic backgrounds. Next, we use LLM Roleplay to simulate dialogues with the same set of personas and goals. Second, we conduct a human evaluation with another pool of participants to assess how well the generated dialogues mimic the collected ones. Specifically, we present participants with two dialogues: one collected from the first study and one simulated using LLM Roleplay, both involving the same persona and the conversation goal.
We then ask the participants to identify which of the dialogues was simulated. We find that LLM Roleplay approximates natural human-chatbot dialogues with a high level of indistinguishability.
@@
Overall, this paper contributes: (i) a novel method to simulate human-chatbot dialogues with an arbitrary set of personas and conversational goals; (ii) a human evaluation confirming our method’s potential to closely resemble real dialogues; (iii) a dataset of goal-oriented human-chatbot and model-chatbot dialogues using our hand-crafted multi-hop goals involving four state-of-the-art LLMs, (iv) an in-depth comparison of open-source and proprietary LLMs in maintaining conversations while embodying specific personas, (v) an open-source implementation of our plug-and-play method, readily applicable for simulating dialogues with any combination of models and chatbot systems across various conversational goals and personas.
@@ 2 Large Language Model Roleplay
At a high level, LLM Roleplay consists of three main steps: (i) Initial conditioning of the inquirer model with persona-specific text alongside the goal description. (ii) Subsequently, the inquirer’s output is provided to the responder model. (iii) Lastly, the output of the responder is returned to the inquirer by asking it to either output a follow-up question or a pre-defined token to terminate the dialogue.
We begin with assembling prompting templates for both of the subjects. We create a system prompt template (SYS_I) for the inquirer LLM. This template prompts the LLM to embody the given persona, provide a prompt to address the designated goal and output the specified termination token if it considers the goal accomplished. As for the responder, we use a default system prompt (SYS_R) to promote it to be a helpful and honest assistant. Additionally, we develop a response forwarder template (INTER_I) for the inquirer. This template prompts the inquirer to assess the conclusiveness of the answer of the responder, determining whether to output a subsequent question or the termination token. We provide the system and forwarder prompt templates in Table 5 in Appendix B. The algorithm for the LLM Roleplay is shown in Algorithm 1.
@@ 3.1 Persona-Specific Dialogue Collection
Subsequently, we generate dialogues utilizing our LLM Roleplay method with a set of three state-of-the-art LLMs and one proprietary conversational agent as inquirers (): llama-2-13B-Chat (Llama-2) Touvron et al. (2023), Mixtral-8x7B-Instruct-v0.1 (Mixtral) Jiang et al. (2024), vicuna-13b-v1.5-16k (Vicuna) Peng et al. (2023), and GPT4 OpenAI et al. (2024). See sample natural and generated dialogues using Mixtral inquirer and Llama-2 responder from Figure 14 to Figure 18 in Appendix E.
In total, the dialogue collection and generation results in the creation of 200 natural human-chatbot () and 800 (with 4 inquirers) simulated () dialogue pairs.
@@
3.2 Impact of Persona-Specific Information
We assess how adding sociodemographic information to the persona description that is provided to the model affects the lexical diversity of the simulated user utterances. We compare three scenarios: (a) using a baseline prompt with no sociodemographic information in the persona description, (b) adding single sociodemographic features, (c) and including all features as described in our method. Table 2 shows the respective Type-Token Ratio (TTR) Zipf (2013) and Distinct-N (dist-n) Li et al. (2016) values quantifying lexical diversity for Mixtral inquirer. We observe that adding single sociodemographic features as well as combined sociodemographic features to the model prompt significantly increases lexical diversity.
| SD Information | TTR | dist-1 | dist-2 |
| None | 0.281 (0.004) | 0.284 (0.003) | 0.625 (0.007) |
| Age | 0.572 (0.010) | 0.576 (0.010) | 0.892 (0.007) |
| Race | 0.582 (0.006) | 0.587 (0.005) | 0.880 (0.002) |
| Gender | 0.411 (0.003) | 0.415 (0.003) | 0.747 (0.002) |
| Education | 0.579 (0.015) | 0.578 (0.016) | 0.869 (0.005) |
| Is Native EN Speaker | 0.394 (0.003) | 0.394 (0.003) | 0.721 (0.005) |
| All | 0.606 (0.019) | 0.605 (0.019) | 0.872 (0.006) |
@@ 3.3 Failure Cases
Since the outputs of the LLMs are free-form, applying deterministic functions to their output proves challenging, resulting in a theoretically infinite number of turns. Therefore, there is a need for a set limit on the number of turns: . Nevertheless, we try to capture the algorithm failure cases to have an automatic assessment of the inquirer LLM failure cases. We expect the inquirer LLM to provide the intended prompt enclosed in double quotes as we explicitly request this in our instructions. See a sample expected output and the failure cases examples in Table 8 in Appendix C. All of the LLMs in our experiments face the following issues.
@@
Toxic content detection.
To proactively address the generation of potentially harmful content, we employ the Llama-2 Guard model Team (2024) to filter out toxic dialogues.
In most cases, GPT4 outperforms other models with lower failure rates: 0.17% for responses without prompts, 0.03% for incoherent responses, and 5.48% for self-replies (Table 1). However, it struggles with providing a single prompt, failing 39.21% of the time. In contrast, Vicuna excels in generating single prompts, achieving a failure rate of 6.08%. Llama-2 receives fewer incoherent outputs, with 0.56%. Mixtral’s performance is intermediate, showing more balanced results across different metrics. Additional plots and detailed data are provided in Appendix B. In these experiments, we detected no unsafe content.
@@ 3.4 Human-Evaluation
To further investigate how various instruction-tuned LLMs behave as inquirers within this setup we analyze the simulated dialogue detection probability, the detection utterance number, and the duration participants spent, using generalized linear mixed models (GLMMs), with the choice of LLM as the independent variable. We additionally include random effects to account for potential confounding effects of individual participants’ detection abilities and users from Section 3.1.
We find significant effects of the choice of the model on the detection probability and the utterance position at which the participants formed their decision. We summarize the results of our statistical analyses in Table 10 in Appendix D using CLD codings (Piepho, 2004) and discuss our key findings in the following.
@@
We fit a binomial model (logit link) to predict the detection probability depending on the model. Concretely, we estimate a GLMM specified by: . We observe a significant effect of the choice of the model on the detection probability (=21.49, ). A post hoc Wald comparison of the contrasts for model types revealed significant differences between Vicuna and all other models (Vicuna being most likely to be detected), and Mixtral and LLama-2 (Mixtral being significantly less likely to be detected). We do not find a significant difference between GPT4 and Mixtral.
@@
Our findings suggest that Mixtral, an open-source LLM, performs better than GPT4 in embodying a specified persona and simulating human-chatbot interactions in terms of detection probabilities. However, in detecting simulated dialogues based on utterance number, GPT4 outperforms, likely due to generating more utterances.
@@ 4 Related Work
Approximation of the true distribution of human-chatbot dialogues is challenging and a significant number of diverse human annotators are needed. Incorporating participants from different sociodemographic profiles addressing a given conversational goal can substantially enhance the richness of the dataset. However, the associated time and resource requirements render the respectively needed large-scale studies infeasible for many researchers. Alternatively, a less resource-intensive and faster method for collecting human-chatbot conversations is to generate them using LLMs Zhang et al. (2024). This method has been utilized in various setups for dialogue generation, such as human-human Adiwardana et al. (2020); Kim et al. (2023); Chen et al. (2023); Li et al. (2023a), teacher-student Macina et al. (2023), and patient-physician Wang et al. (2023).
@@
Large language models (LLMs) have demonstrated distinct behaviors and personas Andreas (2022); Wolf et al. (2024). Andreas (2022) note that LLMs interpret behaviors from text prompts, influencing generated content. Additionally, Wolf et al. (2024) argue that LLMs function as mixture decompositions, where prompts shift component balances, thus triggering persona-specific responses. Furthermore, Beck et al. (2024) investigate the effects of sociodemographic prompts on model responses, finding that while beneficial in some settings, they produce varied effects across models.
Building on these findings, our work introduces a novel, goal-oriented, persona-centric method for generating diverse, multi-turn dialogues. This method simulates dialogues across various combinations of conversational goals, personas, and LLMs, enabling the creation of countless simulated dialogues without theoretical limits.
@@ 5 Discussion and Future Work
In this work, we propose the novel LLM Roleplay method: an automatic, model-agnostic approach for eliciting multi-turn, goal-oriented, persona-based simulated human-chatbot dialogues. We develop and validate our method through two user studies involving 40 participants. Our findings show that up to 44% of these dialogues, where 50% represents perfect indistinguishability, are indistinguishable from real human-chatbot interactions and feature more turns than previous datasets.
Building on this work’s findings, several future research avenues warrant exploration. First, our method can generate user-specific conversational datasets for targeted alignment and domain-adaptation (e.g., using RLHF Christiano et al. (2017)). Second, it can improve dialogue evaluation by providing ample realistic conversational data.
While this paper presents important findings on LLM Roleplay and provides the first evidence that our method can approximate human-chatbot dialogues, the sociodemographic representativeness of our method still needs assessment and advancement. We release our dialogue dataset, method code, and synthesized dialogues to support future research in this promising field.
@@ 6 Conclusion
We present our novel LLM Roleplay method for simulating human-chatbot interaction. In a series of two user studies, we collect real human-chatbot dialogues and demonstrate that LLM Roleplay can generate diverse multi-turn conversations that approximate natural human-chatbot dialogues with a high level of indistinguishability. Our findings highlight the potential of LLMs in simulating human-chatbot interactions to synthesize realistic dialogues that create new opportunities for real-time model evaluation and training data generation for model fine-tuning.